At a certain temperature, the solubility of strontium arsenate, Sr3(AsO4)2, is 0.0510 g/L. What is the Ksp of this salt at this temperature?
Answer #1

0.0510 g = to mol

MW = 540.698

mol = mass/MW = 0.051/540.698 = 0.00009432252

M = 0.00009432252 M

then

Sr3(AsO4)2 --> 3Sr+2 and 2(AsO4)-3

Ksp = [Sr+2]^3[AsO4)-3]^2

then

[Sr+2] = 3S = 3*0.00009432252 = 0.00028296756

[AsO4)-3] = 2S = 2*0.00009432252 = 0.00018864504

Then

Ksp = [Sr+2]^3[AsO4)-3]^2

Ksp = (0.00028296756^3)(0.00018864504^2) = 8.0630*10^-19